# Algebras behind the bispectrality of the Wilson rational functions and their ${}_{4}\phi_{3}$ limits

### v1

@@ 1 Introduction
This paper explores the algebraic structures that underscore the bispectrality of families of biorthogonal rational functions (BRF) of hypergeometric type. The adjective bispectral applies to functions that depend on two sets of variables say, ”coordinates” and ”degrees” and obey two groups of spectral equations one involving operators acting on the coordinates with eigenvalues depending on the degrees and the other equations working in the reciprocal way. This line of research for BRF is in the spirit of the interpretation of the orthogonal polynomials (OP) of the Askey scheme in terms of the Askey-Wilson algebra [1], [2] and its specializations. In the case of the univariate OPs, bispectrality stems from the three-term recurrence relation and the differential or difference equation that the classical polynomials verify. Both equations are of the standard eigenvalue form. The underlying algebra is identified from the relations realized by the operators intervening in these equations and taken in either the coordinate or the degree representation. For finite families of polynomials, this is intimately related to the theory of Leonard pairs [3] which consists in two linear maps that each act on an eigenbasis for the other in a tridiagonal fashion. The corresponding bispectral OPs arise in the overlaps between the two bases of the operators forming the Leonard pair. These studies have much enriched the fruitful interface between orthogonal polynomials, special functions and representation theory and have generated advances in numerous other fields. The present article is part of the efforts aimed at determining how the features of bispectral biorthogonal rational functions can be algebraically encoded and unraveled.
@@ 2 Wilson rational functions
Let us remark that there seems to exist a relation between the RHS of the recurrence and difference GEVP given by, for $n,x$ non-negative integers,
@@
for $n$ and $x$ non-negative integers. This relation has been checked only for different values of $n$ using mathematical software.
@@
${\overline{B}}_{n}(x)={\overline{Z}}_{n-1,n}{(b,c)}\underbrace{(1-aq^{x+n})(cq^{-n+1}/a-q^{-x})}_{\displaystyle=\lambda(q^{x};c)-\lambda(q^{-n}/a;c)}\,,$ (31)
@@
${\overline{X}}_{x+1,x}(c,b)\,{\overline{W}}_{n}(x+1)+{\overline{X}}_{x,x}(c,b)\,{\overline{W}}_{n}(x)+{\overline{X}}_{x-1,x}(c,b)\,{\overline{W}}_{n}(x-1)$
$=$ $\lambda(q^{n};b)\Big{(}{\overline{Z}}_{x+1,x}{(c,b)}\,{\overline{W}}_{n}(x+1)+{\overline{Z}}_{x,x}{(c,b)}\,{\overline{W}}_{n}(x)+{\overline{Z}}_{x-1,x}{(c,b)}\,{\overline{W}}_{n}(x-1)\Big{)}\,.$ (35)
@@ 3 Wilson rational algebra
Let us define the tridiagonal matrices ${\overline{X}}$ and ${\overline{Z}}$ with the entries ${\overline{X}}_{i,j}-(1+cq){\overline{Z}}_{i,j}$ and ${\overline{Z}}_{i,j}$, respectively. These matrices satisfy exactly the relations (39). This statement can be rephrased as $(X,Z)$ and $({\overline{X}},{\overline{Z}})$ provide two different representation of the algebra defined by relations (39).
To take in account of the difference GEVP, one considers their eigenvalues $\lambda(q^{n};b)$. Let also define the diagonal matrix $V$ with non-vanishing entries $V_{n,n}=\lambda(q^{n};b)$ and look for the relations between $V$, $X$ and $Z$. By staigtforward computations, one shows that
$\frac{a}{bc}[Z,V]_{q}+\frac{q}{a}[V,Z]_{q}={\mathbf{d}}_{1}V+{\mathbf{d}}_{2}X+{\mathbf{d}}_{3}{\mathbb{I}}\,,$ (41)
@@
These relations resemble the corresponding relations for the Askey-Wilson (or $q$-Racah) algebra. There are however additional terms in their rhs meaning that the above algebra has a more complicated structure. In particular, in contrast to the $q$-Racah algebra, there are no explicit expressions for the spectra of the operators $X$ and $Z$. In this respect the rational Wilson algebra seems to have a structure that is closer to that of the $q$-Heun-Racah algebra, introduced in [19, 20] and to the one appearing in the study of the Clebsch-Gordan of $sl(3)$ [21, 22].
@@ 4 Limits to $q$-hypergeometric ${}_{4}\phi_{3}$
In this section, limiting cases to ${}_{4}\phi_{3}$ functions presented in [16] are recalled and, for each case, the GEVP are presented.
@@ 4.2 First $q$-Racah rational function
$[{\mathcal{Z}},{\mathcal{X}}]_{q}={\mathfrak{f}}_{1}{\mathcal{Z}}+{\mathfrak{f}}_{2}{\mathcal{X}}\,,$ (51a)
$[{\mathcal{V}},{\mathcal{Z}}]_{q}={\mathfrak{f}}_{2}{\mathcal{V}}+{\mathfrak{f}}_{4}{\mathcal{X}}+{\mathfrak{f}}_{5}{\mathbb{I}}\,,$ (51b)
$[{\mathcal{X}},{\mathcal{V}}]_{q}={\mathfrak{f}}_{1}{\mathcal{V}}+{\mathfrak{f}}_{6}{\mathcal{Z}}+{\mathfrak{f}}_{7}{\mathbb{I}}\,,$ (51c)
@@
The algebra generated by ${\mathcal{X}}$, ${\mathcal{Z}}$ and ${\mathcal{V}}$ subject to the relations (51) stands to be the meta $q$-Racah algebra that will be denoted denoted by $m{\mathfrak{R}}_{q}$. We expect that the meta $q$-Racah algebra generalizes the meta $q$-Hahn algebra and meta Hahn algebra studied in [9]-[10] and completes the set of this type of algebra.
@@
${\overline{W}}_{n}(x)\to{\overline{\mathscr{R}}}^{(1)}_{n}(x)={}_{4}\phi_{3}\left({{q^{-n},\;bq^{n+1},\;q^{-x},\;q/(ade)}\atop{q/d,\;q/e,\;bq^{1-x}}/a}\;\Bigg{|}\;q;q\right)\,,$ (53)
@@
$[\overline{{\mathcal{X}}},\overline{{\mathcal{Z}}}]_{q}={\mathfrak{f}}_{1}\overline{{\mathcal{Z}}}+{\mathfrak{f}}_{2}\overline{{\mathcal{X}}}\,,$ (54a)
$[\overline{{\mathcal{Z}}},\overline{{\mathcal{V}}}]_{q}={\mathfrak{f}}_{2}\overline{{\mathcal{V}}}+{\mathfrak{f}}_{4}\overline{{\mathcal{X}}}+{\mathfrak{f}}_{5}{\mathbb{I}}\,,$ (54b)
$[\overline{{\mathcal{V}}},\overline{{\mathcal{X}}}]_{q}={\mathfrak{f}}_{1}\overline{{\mathcal{V}}}+{\mathfrak{f}}_{6}\overline{{\mathcal{Z}}}+{\mathfrak{f}}_{7}{\mathbb{I}}\,,$ (54c)
@@ 4.3 Other $q$-Racah rational functions
The Wilson rational algebra associated to this functions has the same structure as the one for the Wilson rational function but numerous coefficients vanish: ${\mathfrak{c}}_{1},{\mathfrak{c}}_{3},{\mathfrak{c}}_{4},{\mathfrak{c}}_{7},{\mathbf{e}}_{1},{\mathbf{e}}_{6}\to 0$ and relation (41) simplifies to
$(bcq-a^{2})[Z,V]_{q}=bcq\{V,Z\}-ab(1+q)X-\frac{q}{bcd}\Big{(}(1+q)(1+cd)ab-(1+bd)(a^{2}+qbc)\Big{)}{\mathbb{I}}\,.$ (58)
@@ 5 Concluding remarks
In the framework of a program aimed at extending the Askey scheme to BRF and at providing through the so-called meta algebras a representation theoretic synthesis of the bispectral properties of these functions, we set out to identify the meta $q$-Racah algebra whose definition was still missing. This goal was achieved by analysing the features of the Wilson rational functions from that same perspective and taking limits. This had the remarkable outcome of leading to the definition of the rational Wilson algebra from which the meta $q$-Racah algebra was obtained.
The results presented here open various research avenues many of which having to do with the second route mentioned in the introduction where the starting point are the algebras instead of the functions. First it paves the way for the completion of the unified algebraic treatment of the hypergeometric OPs and BRFs through the construction that can now be done of the representations of the meta $q$-Racah algebras and its $q=1$ version that were identified. Second, from the knowledge of the various examples of meta algebras that have been defined and used, it would be of significant interest to develop the general theory of this class of algebras. Finally, it would be a remarkable accomplishment to show that the rational Wilson algebra encodes the bispectral properties of the eponymous functions by showing that the recurrence coefficients of the GEVP they satisfy can be recovered from the representations of this algebra.

### v2

@@ Abstract
Keywords: Hypergeometric functions; Recurrence relations; Bispectrality; Generalized eigenvalues problem.
@@ 1 Introduction
This paper explores the algebraic structures that underlie the bispectrality of families of biorthogonal rational functions (BRF) of hypergeometric type. The adjective bispectral applies to functions that depend on two sets of variables say, "coordinates" and "degrees" and obey two groups of spectral equations one involving operators acting on the coordinates with eigenvalues depending on the degrees and the other equations working in the reciprocal way. This line of research for BRF is in the spirit of the interpretation of the orthogonal polynomials (OP) of the Askey scheme in terms of the Askey-Wilson algebra [24], [6] and its specializations. In the case of the univariate OPs, bispectrality stems from the three-term recurrence relation and the differential or difference equation that the classical polynomials verify. Both equations are of the standard eigenvalue form. The underlying algebra is identified from the relations realized by the operators intervening in these equations and taken in either the coordinate or the degree representation. For finite families of polynomials, this is intimately related to the theory of Leonard pairs [16] which consists in two linear maps that each act on an eigenbasis for the other in a tridiagonal fashion. The corresponding bispectral OPs arise in the overlaps between the two bases of the operators forming the Leonard pair. These studies have much enriched the fruitful interface between orthogonal polynomials, special functions and representation theory and have generated advances in numerous other fields. The present article is part of the efforts aimed at determining how the features of bispectral biorthogonal rational functions can be algebraically encoded and unraveled.
@@ 2 Wilson rational functions
Let us remark that there exists a relation between the RHS of the recurrence and difference GEVP given by, for $n,x$ non-negative integers,
@@
for $n$ and $x$ non-negative integers. This relation has been proven by using explicit expressions (1) of the Wilson rational functions and looking at the coefficients of, for $k=0,1,\dots n+1$,
$\tau_{k}(x)=\frac{(q^{-x},cq^{x+1};q)_{k}}{(aq^{-x}/c,aq^{x+1};q)_{k}}\,$, (21)
and using the following expression
$Z_{x+1,x}(c)\,\tau_{k}(x+1)+Z_{x,x}(c)\,\tau_{k}(x)+Z_{x-1,x}(c)\,\tau_{k}(x-1)$
$=$ $\omega^{+}(k)\tau_{k+1}(x)+\omega^{0}(k)\tau_{k}(x)+\omega^{-}(k)\tau_{k-1}(x)\,.$ (22)
We do not provide explicit expressions for $\omega^{\pm}(k),\omega^{0}(k)$ here, since they are not useful for the rest of this paper.
@@
${\overline{B}}_{n}(x)={\overline{Z}}_{n-1,n}{(b,c)}\underbrace{(1-aq^{x+n})(cq^{-n+1}/a-q^{-x})}_{\displaystyle=\lambda(q^{x};c)-\lambda(q^{-n}/a;c)}\,$, ($3$)
@@
${\overline{X}}_{x+1,x}(c,b)\,{\overline{W}}_{n+1}(x)+{\overline{X}}_{x,x}(c,b)\,{\overline{W}}_{n}(x)+{\overline{X}}_{x-1,x}(c,b)\,{\overline{W}}_{n-1}(x)$
$=$ $\lambda(q^{n};b)\Big({\overline{Z}}_{x+1,x}{(c,b)}\,{\overline{W}}_{n+1}(x)+{\overline{Z}}_{x,x}{(c,b)}\,{\overline{W}}_{n}(x)+{\overline{Z}}_{x-1,x}{(c,b)}\,{\overline{W}}_{n-1}(x)\Big)\,.$ (37)
@@ 3 Wilson rational algebra
Let us define the tridiagonal matrices $\overline{X}$} and $\overline{Z}$} with the entries ${\overline{X}}_{i,j}$ and ${\overline{Z}}_{i,j}$, respectively. These matrices satisfy exactly the relations (41). This statement can be rephrased as $(X,Z)$ and $({\overline{X}},{\overline{Z}})$ provide two different representation of the algebra defined by relations (41).
To take in account of the difference GEVP, one considers their eigenvalues $\lambda(q^{n};b)$. Let also define the diagonal matrix $V$ with non-vanishing entries $V_{n,n}=\lambda(q^{n};b)$ and look for the relations between $V$, $X$ and $Z$. By straightforward computations, one shows that
$\frac{a}{bc}[V,Z]_{q}+\frac{q}{a}[Z,V]_{q}={\mathbf{d}}_{1}V+{\mathbf{d}}_{2}X+{\mathbf{d}}_{3}{\mathbb{I}}\,$, (43)
@@
These relations resemble the corresponding relations for the Askey-Wilson (or $q$-Racah) algebra. There are however additional terms in their rhs meaning that the above algebra has a more complicated structure. In particular, in contrast to the $q$-Racah algebra, there are no explicit expressions for the spectra of the operators $X$ and $Z$. In this respect the rational Wilson algebra seems to have a structure that is closer to that of the $q$-Heun-Racah algebra, introduced in [2, 3] and to the one appearing in the study of the Clebsch-Gordan problem of $sl(3)$ [7, 8].
@@ 4 Limits to $q$-hypergeometric $_{4}\phi_{3}$
In this section, limiting cases to $_{4}\phi_{3}$ functions presented in [10] are recalled and, for each case, the GEVP are presented. Let us mention [21] where these limits are classified and studied in detail.
@@ 4.2 First $q$-Racah rational function
$[{\mathcal{X}},{\mathcal{Z}}]_{q}={\mathfrak{f}}_{1}{\mathcal{Z}}+{\mathfrak{f}}_{2}{\mathcal{X}}\,$, (53a)
$[{\mathcal{Z}},{\mathcal{V}}]_{q}={\mathfrak{f}}_{2}{\mathcal{V}}+{\mathfrak{f}}_{4}{\mathcal{X}}+{\mathfrak{f}}_{5}{\mathbb{I}}\,$, (53b)
$[{\mathcal{V}},{\mathcal{X}}]_{q}={\mathfrak{f}}_{1}{\mathcal{V}}+{\mathfrak{f}}_{6}{\mathcal{Z}}+{\mathfrak{f}}_{7}{\mathbb{I}}\,$, (53c)
@@
The algebra generated by $\mathcal{X}$}, $\mathcal{Z}$} and $\mathcal{V}$} subject to the relations (53) stands to be the meta $q$-Racah algebra that will be denoted by $m{\mathfrak{R}}_{q}$. We expect that the meta $q$-Racah algebra generalizes the meta $q$-Hahn algebra and meta Hahn algebra studied in [20]-[4] and completes the set of this type of algebra..
@@
${\overline{W}}_{n}(x)\to{\overline{\mathscr{R}}}^{(1)}_{n}(x)={}_{4}\phi_{3}\left({{q^{-n},\;bq^{n+1},\;q^{-x},\;q/(ade)}\atop{q/d,\;q/e,\;bq^{1-x}}/a}\;\Bigg|\;q;q\right)\,$, ($5$)
@@
$[\overline{{\mathcal{Z}}},\overline{{\mathcal{X}}}]_{q}={\mathfrak{f}}_{1}\overline{{\mathcal{Z}}}+{\mathfrak{f}}_{2}\overline{{\mathcal{X}}}\,$, (56a)
$[\overline{{\mathcal{X}}},\overline{{\mathcal{V}}}]_{q}={\mathfrak{f}}_{2}\overline{{\mathcal{V}}}+{\mathfrak{f}}_{4}\overline{{\mathcal{X}}}+{\mathfrak{f}}_{5}{\mathbb{I}}\,$, (56b)
$[\overline{{\mathcal{V}}},\overline{{\mathcal{Z}}}]_{q}={\mathfrak{f}}_{1}\overline{{\mathcal{V}}}+{\mathfrak{f}}_{6}\overline{{\mathcal{Z}}}+{\mathfrak{f}}_{7}{\mathbb{I}}\,$, (56c)
@@ 4.3 Other $q$-Racah rational functions
The Wilson rational algebra associated to this functions has the same structure as the one for the Wilson rational function but numerous coefficients vanish: ${\mathfrak{c}}_{2},{\mathfrak{c}}_{4},{\mathfrak{c}}_{5},{\mathfrak{c}}_{10},{\mathbf{d}}_{2},{\mathbf{e}}_{1},{\mathbf{e}}_{7}\to 0$.
@@ 5 Concluding remarks
In the framework of a program aimed at extending the Askey scheme to BRF and at providing through the so-called meta algebras a representation theoretic synthesis of the bispectral properties of these functions, we set out to identify the meta $q$-Racah algebra whose definition was still missing. This goal was achieved by analysing the features of the Wilson rational functions from that same perspective and taking limits. This had the remarkable outcome of leading to the definition of the rational Wilson algebra from which the meta qqRacah algebra was obtained.
The results presented here open various research avenues many of which having to do with the second route mentioned in the introduction where the starting point are the algebras instead of the functions. First it paves the way for the completion of the unified algebraic treatment of the hypergeometric OPs and BRFs through the construction that can now be done of the representations of the $q$-meta Racah algebras and its $q=1$ version that were identified. Second, from the knowledge of the various examples of meta algebras that have been defined and used, it would be of significant interest to develop the general theory of this class of algebras. Finally, it would be a remarkable accomplishment to show that the rational Wilson algebra encodes the bispectral properties of the eponymous functions by showing that the recurrence coefficients of the GEVP they satisfy can be recovered from the representations of this algebra.